# Uncovering the Functional Roles of Nonlinearity in Memory

### v1

@@ 4 Experimental Results
To probe how recurrent nonlinearity shapes memory, we selected a set of tasks spanning those that can be solved by linear dynamics and those that require nonlinear computation, allowing us to highlight distinct linear and nonlinear memory mechanisms (Sect. 3) engaged by different computational demands.
4.1 Nonlinearity Reshapes Slow Modes in Temporal Integration
We begin with a class of tasks that require gradual accumulation of information over time to support a final decision. At each time step, the model receives a new input $\bm{x}_{t}$, and after a sequence of $T$ steps, it must predict a global output $\bm{y}$, such as a class label. We test this setup across diverse input modalities, including images, audio, and text. The setup mirrors the structure of drift diffusion processes in psychology [54], in which evidence is integrated over time until a decision threshold is reached. Analogously, the AL-RNN’s latent state $\bm{z}_{t}$ integrates incoming representations $\bm{s}_{t}$ over time, and a prediction $\hat{\bm{y}}$ (e.g., a class label) is made from the final state $\bm{z}_{T}$.
Sentiment Classification on IMDb
We first analyzed binary sentiment classification using the IMDb dataset [39]. Input sequences consist of tokenized movie reviews, which were standardized by truncation or padding to a sequence length of $128$ during training and testing. We mapped raw tokens $\bm{x}_{t}$ to latent inputs $\bm{s}_{t}$ by combining pretrained GloVe vectors [51] with a trainable nonlinear embedding layer. Fig. 1a illustrates the overall scheme. As depicted in Fig. 1b, varying the number of nonlinear units minimally affects performance, suggesting - as may be expected - a relatively simple linear temporal integration mechanism. During a negative review (Fig. 1d) large updates to the latent state occur almost exclusively when highly valenced keywords such as worst or terrible are encountered, confirming that memory integration is dominated by sparse lexical cues already emphasized by the input embedding. The distribution of nonlinear bitcodes is correspondingly degenerate (Fig. 1c): a handful of codes account for almost all examples.
Dynamically, temporal integration is driven by a dominant slow mode, reflected in a single eigenvalue near 1, with all others substantially smaller, indicating faster-decaying transients (Fig. 1e). This configuration enables the system to maintain and integrate information over time primarily along the slowest dynamical mode. Supporting this, the eigenvector associated with the slow mode is almost perfectly aligned (cosine similarity > 0.999) with the first PC of the latent trajectories, which alone accounts for $98\%$ of the total variance—indicating that the system’s activity is effectively constrained to a one-dimensional (sentiment-related) integration axis. Similar slow-mode or line attractor based integration mechanisms have been implicated in memory and evidence accumulation tasks across both artificial and biological systems [58, 40].
Figure 1: a: Schematic of the processing pipeline. b: Test accuracy as a function of nonlinear units, $P$. c: Bitcode frequencies revealing clustering of latent states. d: Trajectory through latent space showing sentiment-relevant keywords guiding classification. e: The dominant first PC clearly separates positive and negative reviews. The eigenvalue spectrum (right) has a maximum value close to one whose eigenvector aligns closely with the first PC.
Sequential MNIST (sMNIST)
@@ 4.1 Nonlinearity Reshapes Slow Modes in Temporal Integration
Figure 2: a: Accuracy (y-axis) vs. number of nonlinear units (x-axis; mean $\pm$ std over 10 seeds). b: Final latent state projection onto the first 3 PCs show that nonlinearity partitions latent space by class.
In the sMNIST task, $28\times 28$ digit images are flattened into a sequential time series of length 784, with one pixel input per time step. Classification is based solely on the final latent state. To preserve sequential structure while enabling feature extraction, we apply a local 1D convolutional encoder before passing the sequence to the AL-RNN (Appx. A.1.3).
As for IMDb, we observe a similar slow accumulation of evidence along the first PCs, but now implementing a more complex, higher-dimensional data manifold (see Figs. 2&12). Accordingly, linear models already perform well on this task, with maximum eigenvalues slightly above 1 ($\lambda_{max}\approx 1.003\pm 0.003$, Fig. 12). However, as shown in Fig. 2a, increasing the number of nonlinear units still improves test accuracy up to an intermediate point, after which performance declines again for fully nonlinear models. Fig.2b shows final latent states projected onto the first three PCs, revealing how nonlinearity enhances class separation by partitioning the latent manifold, achieving up to $98.6\%$ test accuracy. Bitcode analysis provides a clear mechanism by which nonlinearity increases performance: each class aligns with its own set of closely neighboring bitcodes. For instance, for $P=50$, the average variation in the bitcode within each class was just $0.06\pm 0.02$ (see Fig. 13), meaning that on average $48$ out of $50$ PWL units had the same sign at the final state. This allows the AL-RNN to implement class-specific linear integration dynamics, resulting in well-separated clusters in latent space (Figs. 2b & 14). These results support the view that nonlinearity in this setting acts not as a core computational mechanism, but as a tool for improving expressivity. However, with full nonlinearity ($P=100$), this separating effect leads to more imbalanced variance distributions across classes (Fig. 14), degrading performance. We observed qualitatively identical behavior in an audio-based classification task (Speech Commands), reported in Appx. A.4.1.
@@ 4.2 Minimal Nonlinearity Stabilizes Transients For Structured Memory Recall
Next, to evaluate the AL-RNN’s ability to sustain stable internal representations over time, we tested it on the classic copy task (Fig. 3a). During the encoding phase, the network receives a random sequence of $N_{sym}$ distinct symbols, presented one per time step. This is followed by a delay of length $D$ time steps with no input, and then a “cue” signal indicates that the network should reproduce the original sequence over the next $N_{seq}$ time steps (the “recall phase”). Performance is measured as the number of sequences recalled correctly. This task is particularly interesting from a neuroscience perspective, as it parallels experimental paradigms used to study working memory in prefrontal cortex, where stable task-relevant information must be maintained over variable delay intervals despite complex and heterogeneous underlying dynamics [47, 53].
Figure 3: a: Top: Structure of the copy task and two example trajectories for $P=1$. b: Top: Activity of the PWL unit ($P=1$) for the same two input sequences. The latent activity follows a complex limit cycle primarily located in one linear subregion (PWL unit negative), which switches to the second subregion during decoding (PWL unit positive). Bottom: Autonomous activity of the AL-RNN in the absence of inputs encodes a 100-cycle, with its transient located only in one subregion (PWL unit negative). c: Symbol-wise recall accuracy (mean±$\pm$ std over 10 seeds) vs. number of PWL unit $P$. d: Histogram of binary “bitcodes” during recall for $P=10$, concentrated on a small subset out of $10242^{10}=$ 1024 possible bitcodes. e: Explained variance ratio of PCs of latent network activity (for $P=1$) indicates relatively high-dimensional, complex dynamics.
@@
In this configuration $(N_{sym}=4,N_{seq}=8$ and $D=200)$, we first find that purely linear models perform above chance, suggesting they can partially solve the task. This aligns with theory: linear systems can support marginally stable oscillations when eigenvalues lie on the unit circle, enabling long recurrence times that could, in principle, encode sequence and timing (see Fig. 15). However, such solutions are highly fragile to noise and perturbations [61]. Minimal nonlinearity therefore greatly improves performance and robustness. A single nonlinear unit ($P=1$) often yields perfect recall, with performance declining for increased nonlinear units and being worst for fully nonlinear models (Fig. 3c). Crucially, we find that even fully nonlinear models often only use a small subset of available linear subregions (Fig. 3d, Fig. 16 for a full statistical analysis).
@@
To understand the underlying computational mechanism, we analyzed a representative AL-RNN with $P=1$ that perfectly solves the task. We find that the PWL unit leverages two distinct linear subregions to separate encoding/decoding vs. storage dynamics. During the encoding and decoding phases, the network remains confined to the primary linear subregion optimized for integrating and recalling symbols (Fig. 3b, top). In contrast, during the delay phase, the network transitions to the second linear subregion, where it implements a complex transient cycle that is structured according to the encoded sequence.
The autonomous dynamics of the model establish a global k-cycle with a precise period of 100 time steps (exactly half the delay phase duration), as confirmed by a neutral maximum Lyapunov exponent ($\lambda_{max}\approx 0$, Fig. 3b, bottom). While the initial transient dynamics are constrained entirely to the lower subregion, the fully developed k-cycle spans both subregions, oscillating between them in a structured manner. This cycle is stabilized through the interplay of two distinct spectral properties: one with a contracting virtual fixed point ($\lambda_{max}\approx 0.992$) and the other around a diverging fixed point ($\lambda_{max}\approx 1.005$) which drives the initial transient dynamics. This structured alternation between contraction and divergence, causing periodic transients, allows the model to robustly maintain memory over long horizons without external input (Fig. 3b).
@@ 4.3 Nonlinearity Enables Gating in the Addition Task
Figure 4: a: Performance as a function of units $P$ (mean±$\pm$ std over 10 seeds). b: Latent trajectories in PWL space with $P=2$ nonlinear units. The second linear subregion is selectively activated only at the two masked time points ("input"), indicated by sharp transitions across quadrant boundaries. Outside these events, the network remains in a single linear regime, leading to smooth integration of the cumulative sum. c: Time course of masked inputs (red), network outputs (grey/black), and target values (blue) for two example trials, where the input occurs either early (light grey) or late (black) in the sequence. In both cases, the trajectory initially follows a linear drift prior to input, then transitions sharply to an elevated integration path with the same slope—offset to reach the target after 100 steps.
@@
Next, we consider the addition problem [28], an established test of selective memory. The input consists of a continuous stream of random numbers, along with a sparse binary mask indicating two marked time points. The goal is to output the sum of the two marked times, ignoring all other inputs. Crucially, the mask is not directly fed into a separate gating mechanism, but the AL-RNN must internally learn to accumulate inputs only at the relevant times. This requires the network to integrate incoming values differently depending on the input—something linear RNNs are incapable of.
@@
Our pipeline accurately identifies that, without nonlinearity, the network fails on this task (Fig. 4a). In contrast, with just one nonlinear unit ($P=1$) the AL-RNN achieves orders-of-magnitude lower error. Fig. 4b illustrates that nonlinearity enables the emergence of an internal gating mechanism, where the AL-RNN learns to selectively "open" and "close" an internal integration pathway conditional on the input mask. To sum the inputs, the network leverages a slow linear drift that accumulates over the full sequence length (Fig. 4c), combined with selectively adding the gated inputs. This mechanism cleanly separates the functional contributions of linear and nonlinear units and aligns with theoretically derived optimal PWL solutions for this task previously described in the literature [46, 57]. Interestingly, as before, even with increased nonlinearity (Fig. 17), the network preserves this strategy, distributing linear integration and gating across two distinct small sets of subregions. However, fully nonlinear models fail to capture this simplicity effectively: their expressivity leads them to "overfit" the integration process, fragmenting the linear accumulation mechanism (Fig. 17). This "over-approximation" of a fundamentally simple process introduces instability, resulting in degraded performance (Fig. 4a).
4.4 Nonlinearity Enables Internal Task Switching Through Context-Dependent Routing
We next assess the network’s ability to switch between tasks using a context-dependent integration task (Fig. 5a), a design inspired by behavioral paradigms in neuroscience for studying flexible decision-making [42]. Each trial begins with a one-hot context cue, followed by a sequence of scalar inputs ($T_{seq}=100$). This task requires the network to switch between two internal decision policies based on a contextual cue. Purely linear AL-RNNs cannot solve this task, and plateau at $50\%$ accuracy (Fig. 5b), independent of how many linear units are provided. In contrast, the task is almost perfectly solvable with just one linear and one PWL unit, achieving up to $96\%$ accuracy.
4.5 Nonlinearity Enables Task Switching in a Stimulus Selection Task in Joint Task–Neural Modeling
Figure 6: a: Schematic of the contextual stimulus selection task used in the PFC-1 dataset [55]. In each trial, animals responded to auditory stimuli based on either the spatial origin (left/right) or the pitch (high/low), depending on the current task rule. b: Latent trajectories from a task-trained 2D model for two identical stimuli under different rules, illustrating context-dependent separation at decision time by the subregion boundary (black horizontal line). c: Task performance across models of increasing nonlinearity (P) and different latent dimensions (M) (left) and misclassification counts by stimulus and block (right). d: Mean firing rate MSE with and without inputs. e: Joint training performance when modeling both task and spike data: test accuracy (left), mean spike rate MSE (middle left), and coefficient of variation (CV) MSE (middle right) and mean spike correlation (right) as a function of the number of nonlinear units. f: Mean activation over time of a representative nonlinear unit that differentiates rules and modulates stimulus-driven activity conditionally.
To test our approach on empirical neuroscientific data, we evaluated it on neural population recordings from the CRCNS PFC-1 dataset [55], which captures single-unit activity in rodent auditory cortex during a flexible stimulus selection task. In this paradigm, animals are required to alternate between two rules: selecting the appropriate response based either on the spatial origin of the sound or on its pitch (Fig. 6a). The two stimulus dimensions are presented simultaneously but independently, with either a low or high warble (pitch cue) played on both sides, and a pure tone (side cue) played either on the left or right. Depending on the current rule, the animal must attend to either the pitch (low = "go", high = "no-go") or the side (left = "go", right = "no-go") of the auditory input. This setup creates contextual ambiguity: for example, an identical stimulus combination of a high-pitched warble and a tone on the left requires different responses ("go" under the spatial rule, "no-go" under the pitch rule). Hence the task studies linear and nonlinear neural computations in context-dependent behavior, with sparse, variable recordings posing a realistic challenge. Preprocessing steps and selection criteria are detailed in the Appx. A.3. To ensure a consistent task representation, we restricted training and evaluation to the $723$ trials in which the animal made the correct choice.
Task-training the AL-RNN
We first trained an AL-RNN to solve the task based purely on structured temporal input. Each trial mimicked one of the 723 empirically observed trials, beginning with a brief context cue indicating the active rule (spatial or pitch), presented for 100 ms (two time steps), followed by a one-second delay. Then, an auditory stimulus was presented (one-hot encoded for pitch and side), and the model was required to produce a binary decision (go/no-go), which was read out linearly from the latent state at the empirically observed decision time $\bm{z}_{t_{dec}}$ (see Appendix A.3). Despite the task’s contextual nature, we found that a 2D model was already often sufficient to linearly separate identical stimuli based on their preceding rule cue, and linear 3D models often solved the task perfectly. Systematic evaluation confirmed that increasing nonlinearity beyond this dimension did not yield substantial gains in task accuracy alone (Fig. 6c, left), and that errors were largely restricted to contextually ambiguous stimuli (e.g., Right–Low under the spatial rule; Fig. 6c, right).
Task Information Improves Spike Reconstruction
We next used the same AL-RNN to jointly model both the animal’s task behavior and the neural spike trains. That is, the network received identical inputs as in the task-only training—context cue, delay, and stimulus—and was required not only to produce the correct decision at the appropriate time, but also to reproduce the observed spiking activity across neurons from the same latent states of the AL-RNN (Eq. 1). To achieve this, we trained the AL-RNN within the Multimodal Teacher Forcing (MTF) framework [10], which allows direct training on non-Gaussian spike data via a Poisson decoder. The decoder is hierarchically conditioned on a 5-dimensional, trial-specific feature vector [11], enabling it to capture slow drift in spike statistics while preserving a shared task representation in the AL-RNN. See Appx.A.1.4 for training details and Fig.20 for an analysis of the learned feature vectors.
We found that including task information markedly improved spike reconstruction (Fig. 6d). Providing the model with the rule cue and stimulus enabled it to better predict the neural data, suggesting a strong alignment between latent decision dynamics and observed neural variability. Quantitatively, spike rates and coefficients of variation were well captured (see Fig. 19a). To evaluate the model’s temporal reconstruction, we correlated recorded and generated spike trains, yielding a mean correlation of $r=0.359\pm 0.002$. This approached the correlation between separate Poisson samples drawn from the same latent trajectory ($r=0.371\pm 0.057$), suggesting an upper bound imposed by the intrinsic variability of the Poisson process.
Nonlinearity Supports Integration Under Dual Constraint
In the joint setting, nonlinearity began to play a more significant role. As shown in Fig. 6e, increasing the number of nonlinear units improves both task accuracy and spike reconstruction ($r\approx-0.37,p=0.005$, with lower firing rate MSE and higher task accuracy being better). In contrast, linear models plateau around $87\%$ accuracy and exhibit an almost identical error pattern: they consistently respond “go” to all trials with left or low cues, regardless of the active rule, thus failing to resolve the contextual ambiguity (see Fig. 6c, right). In many such nonlinear models, we observed the emergence of single units that effectively gated the task rule. These units became active in one rule context and inactive in the other, with stimulus integration occurring only within the active subregion, providing an interpretable mechanism for context-dependent switching (Fig. 6f). Quantitatively, we examined the PWL subregion active at the time the input is first presented (after one second) and compared which subregions were used in blocks 2 and 4 (the two conditions where both spatial and pitch cues are present but different rules are active). In successful models ($>99\%$ accuracy), trials from these blocks were routed into almost entirely distinct subregions (on average, $97\%$ of trials fell into block-specific linear subregions), while in unsuccessful models this separation was much less pronounced (only $32\%$). Accuracy was strongly correlated with this separation score ($r=0.66$). This shows that successful models leverage nonlinearity to assign different task rules to distinct linear regimes, allowing them to react differently to identical stimuli depending on context. This mechanism is illustrated for a successful task-trained 2D model in Fig. 6b.
Our results underscore an interesting tension in jointly modeling task behavior and neural activity: while the task is linearly solvable in 2D when trained on the task alone, introducing the sparse, noisy and variable cortical spike data, imposes additional complexity that benefits from the additional expressivity provided by nonlinear models.

### v2

@@ 4 Experimental Results
To probe how recurrent nonlinearity shapes memory, we selected a set of tasks spanning those that can be solved by linear dynamics and those that require nonlinear computation, allowing us to highlight distinct linear and nonlinear memory mechanisms (Sect. 3) engaged by different computational demands. In addition, to assess the robustness of the framework under more challenging conditions, we apply it to a multi-task setting and to an empirical neuroscientific dataset.
Our primary goal is interpretability rather than performance optimization, making direct architectural comparisons less straightforward. However, to contextualize AL-RNN performance, we included standard gated RNN baselines (LSTM, GRU) on the performance-oriented benchmarks (sMNIST, Speech Commands, Copy, Addition), and investigated the influence of sparse nonlinearity within S4 and a Linear Transformer on the Copy and Addition task (see Fig. 27 and Table 2). We also compared AL-RNNs with different activation functions (ReLU, GeLU, tanh, hardtanh), finding qualitatively similar results (see Table 3), though some systematic differences emerge, discussed further in Sect. 4.3.
4.1 Modality-Specific Encoding Enables Simple Linear Integration
We begin with a class of tasks that require gradual accumulation of information over time to support a final decision. At each time step, the AL-RNN’s latent state $\bm{z}_{t}$ integrates incoming representations $\bm{s}_{t}$ over time, and the class label $\hat{\bm{y}}$ is predicted from the final state $\bm{z}_{T}$. We test this setup across images, audio, and text, employing domain-appropriate encoders to map raw inputs to latent representations: pretrained GloVe embeddings (Pennington et al., 2014) for text, 1D convolutional layers for sequential pixels, and spectral features for audio (see Appx. A.1.3). While these encoders necessarily differ in their treatment of modality-specific structure, they serve the common purpose of transforming heterogeneous inputs into a shared representational space amenable to temporal integration. For instance, the GloVe embeddings provide semantically meaningful representations where valenced words (e.g., "good" vs. "bad") are already separable, putting focus on the temporal integration of information rather than learning semantic structure from scratch. By constraining the recurrent dynamics to operate on these normalized representations and requiring classification through a linear readout, we enforce that the main computational mechanism, the temporal integration itself, is expressed explicitly in the AL-RNN’s dynamics, revealing that despite the diversity of input modalities, the underlying task solution converges to the same dynamical principle across all three domains.
@@ 4.1 Modality-Specific Encoding Enables Simple Linear Integration
Figure 2: Top row: Sentiment classification on IMDb reviews. a: Test accuracy (y-axis) as a function of nonlinear units, $P$. b: Trajectory through latent space showing sentiment-relevant keywords guiding classification. c: The dominant first PC clearly separates positive and negative reviews. d: Bitcode frequencies are highly concentrated on a small subset of linear subregions. Bottom row: Digit classification on sequential MNIST. e: Test accuracy (y-axis) as a function of nonlinear units (x-axis; mean $\pm$ std over 10 seeds). f: Final latent state projection onto the first 3 PCs show that nonlinearity partitions latent space by class.
For binary sentiment classification using the IMDb dataset (Maas et al., 2011), the integration mechanism is remarkably simple, and increasing nonlinearity has minimal effect on performance (Fig. 2a; Welch’s $t$-test comparing $P=0$ vs $P=100$: t=1.34t=-1.34, $p=0.20$). During a negative review (Fig. 2b), large updates to the latent state occur almost exclusively when highly valenced keywords such as worst or terrible are encountered, suggesting that memory integration is dominated by sparse lexical cues already emphasized by the input embedding. Accordingly, the dynamics are dominated by a single slow mode: an approximate line attractor with one eigenvalue near 1 and all others substantially smaller  [^1] We use the term ”line attractor” in a functional sense: the dynamics exhibit very slow integration along a dominant mode (eigenvalue $\approx 1$) that preserves information over task-relevant timescales. Strictly speaking, true line attractors require a continuous manifold of neutrally stable fixed points, while our linear models instead have a single fixed point with one very slow decay mode.. The eigenvector associated with the slow mode is almost perfectly aligned (cosine similarity > 0.999), which alone accounts for $98\%$ of the total variance, indicating that the system’s activity is effectively constrained to a one-dimensional (sentiment-related) integration axis (Fig. 2c), mirroring prior observations in the literature (Maheswaranathan et al., 2019). The distribution of nonlinear bitcodes is correspondingly degenerate (Fig. 2d): a handful of codes account for almost all examples.
We observed similar slow-mode integration mechanisms in image classification (Sequential MNIST= sMNIST) and audio classification (Speech Commands), detailed in Appx. A.4.2. While the overall dynamical mechanism is very similar, in both cases, the data manifold is higher-dimensional and requires more complex latent representations. Here, sparse nonlinearity provides a modest but significant improvement (Fig. 2e for sMNIST, Fig. 13 for Speech Commands).
For sMNIST, bitcode analysis reveals a clear mechanism by which nonlinearity affects performance: each of the 10 digit classes aligns with its own set of closely neighboring bitcodes, enabling class-specific linear integration dynamics (Fig. 10). While all classes are processed through slow modes with eigenvalues near 1 (Fig. 9), small differences in these eigenvalues, arising from distinct linear subregions, cause trajectories to diverge into more well-separated clusters (Fig. 2f). Performance peaks at intermediate nonlinearity and declines for fully nonlinear models, which overdisperse the data manifold (Fig. 12). The fact that this same slow-mode integration structure emerges across text, images, and audio demonstrates a universal computational motif for evidence accumulation tasks, consistent with similar mechanisms observed in biological systems (Seung, 1996, Ratcliff and McKoon, 2008, Maheswaranathan et al., 2019).
@@ 4.2 Minimal Nonlinearity Stabilizes Transients For Structured Memory Recall
Next, to evaluate the AL-RNN’s ability to sustain stable internal representations over time, we tested it on the classic copy task (Fig. 3a). During the encoding phase, the network receives a random sequence of $N_{sym}$ distinct symbols, presented one per time step. This is followed by a delay of length $D$ time steps with no input, and then a “cue” signal indicates that the network should reproduce the original sequence over the next $N_{seq}$ time steps (the “recall phase”). Performance is measured as the number of sequences recalled correctly. This task is particularly interesting from a neuroscience perspective, as it parallels experimental paradigms used to study working memory in prefrontal cortex, where stable task-relevant information must be maintained over delay intervals despite complex and heterogeneous underlying dynamics (Murray et al., 2017, Rajan et al., 2016).
[…]
@@
In this configuration $(N_{sym}=4,N_{seq}=8$ and $D=200)$, we first find that purely linear models perform above chance, suggesting they can partially solve the task. This aligns with theory: linear systems can support marginally stable cycles when eigenvalues lie on the unit circle, enabling long recurrence times that could, in principle, encode sequence and timing. However, such solutions are highly fragile to noise and perturbations (Strogatz, 2018), confirmed by the observation that despite extensive hyperparameter tuning, linear models fail to solve this task with high accuracy (Fig. 3c). Minimal nonlinearity therefore greatly improves performance and robustness. A single nonlinear unit ($P=1$) often yields perfect recall, with performance declining for increased nonlinear units and being worst for fully nonlinear models (Fig. 3c).
@@
To understand the underlying computational mechanism underlying successful solutions, we analyzed a representative AL-RNN with $P=1$ that perfectly solves the task, i.e. decodes all encoded symbols with $100\%$ accuracy on the test set. We find that the PWL unit leverages two distinct linear subregions to separate encoding/decoding vs. storage dynamics. During the encoding and decoding phases, the network remains confined to the primary linear subregion optimized for integrating and recalling symbols (Fig. 3b, top). In contrast, during the delay phase, the network transitions to the second linear subregion, where it implements a complex transient cycle that stores the encoded sequence. Because the encoded sequence uniquely determines the entry point into the cycle’s transient, the network preserves this phase information throughout the delay, allowing the sequence to be decoded from the corresponding exit point at recall.
Crucially, we find that sparsely nonlinear models often only use a small subset of available linear subregions (Fig. 3d, Fig. 16 for a full statistical analysis). Fig. 17a illustrates why fully nonlinear models fail to converge to good solutions: their loss curves are highly irregular and feature frequent bifurcations, directly linked to high gradient norms throughout training. For trained models, solutions are spread out across hundreds of subregions (bitcodes), failing to learn simple task solutions (Fig. 18a). For results on the variable delay task, where the recall cue occurs after a random interval, see Sect. A.4.3.
@@ 4.3 Nonlinearity Enables Gating in the Addition Task
[…]
@@
Next, we consider the addition problem (Hochreiter and Schmidhuber, 1997), an established test of selective memory. The input consists of a continuous stream of random numbers, along with a sparse binary mask indicating two marked time points. The goal is to output the sum of the two marked times, ignoring all other inputs. Crucially, the mask is not directly fed into a separate gating mechanism, but the AL-RNN must internally learn to accumulate inputs only at the relevant times. This requires the network to integrate incoming values differently depending on the input. This task is impossible to solve by a linear AL-RNN (see Appx. A.4.1 for a formal proof), as accurately identified by our pipeline (Fig. 4a). In contrast, with just a single nonlinear unit ($P=1$) the AL-RNN achieves orders-of-magnitude lower error.
@@
Fig. 4b illustrates that nonlinearity enables the emergence of an internal gating mechanism, where the AL-RNN learns to selectively "open" and "close" an internal integration pathway conditional on the input mask. To sum the inputs, the network leverages a slow linear drift that accumulates over the full sequence length (Fig. 4c), while nonlinear units rapidly reposition the state onto the appropriate integration trajectory when gated inputs occur. This mechanism cleanly separates the functional contributions of linear and nonlinear units and aligns with theoretically derived optimal PWL solutions for this task previously described in the literature (Monfared and Durstewitz, 2020, Schmidt et al., 2021). Interestingly, as before, even with increased nonlinearity, the network preserves this strategy (Fig. 14a), distributing linear integration and gating across two distinct small sets of subregions. While fully nonlinear models can outperform sparsely nonlinear models (Fig. 18b), some fully nonlinear models fail to capture this simplicity of the task mechanism effectively, fragmenting the linear accumulation mechanism (Fig. 18b, middle vs. right), an effect that is conserved across training and test set (Fig. 14c). In line with this, we find that fully nonlinear models frequently “linearize” units by shifting their mean activity far from zero, effectively enforcing linear dynamics but at the cost of more difficult optimization (Fig. 14b).
Interestingly, all tested nonlinearities showed a similar performance pattern: marked improvement at $P=1$, modest gains with sparse nonlinearity, stable plateaus at intermediate $P$, and worse performance at full $P$ (Table 3). What varied was how cleanly interpretable mechanisms emerged. The addition task requires combining two linear processes: slow drift accumulating time, and selective integration of marked inputs. PWL activations like ReLU naturally separate these into distinct linear subregions, whereas $\tanh$ must approximate the separation by shifting activity between near-linear and saturated regimes. This softer partition is harder to learn and yields roughly an order of magnitude higher MSE, highlighting the inductive bias of PWL units for switching-linear tasks.
4.4 Nonlinearity Facilitates Routing in Contextual Integration
Beyond simple integration, we examined context-dependent tasks where the model must respond differently to identical stimuli based on preceding rule cues. We tested this in two settings: a simple benchmark requiring conditional integration of a stimulus sequence, and a neuroscientific task involving joint modeling of both behavioral decisions and neural population activity from rodent auditory cortex (PFC-1 dataset, Rodgers and DeWeese (2014)). For the latter, we jointly trained the AL-RNN to both solve the task and reconstruct neural population activity from the same latent states.
In both settings, linear models fail completely, plateauing at chance performance, while introducing even a single nonlinear unit enables near-perfect accuracy (Appx. A.4.5& A.4.4). In both cases, the central mechanism was straightforward to interpret: nonlinearity routes different task rules into distinct linear subregions, implementing context-dependent gating that linear dynamics fundamentally cannot achieve. While the overall task exhibits nonlinear dependencies, requiring opposite responses to identical inputs depending on context, the piecewise structure allows each subregion to implement its own linear solution to its respective subtask. Building on this insight that nonlinearity enables computational routing, we next explore how sparse nonlinearity structures computational reuse across multiple related tasks.
4.5 Shared Nonlinear Structure in a Multi-Task Paradigm
Figure 5: AL-RNN training on a multi-task paradigm. (a) Test accuracy (mean across 5 models per setting) across 11 cognitive tasks as a function of $P$ (PWL units), for different training set sizes. For limited data (20 samples per task), sparse nonlinearity ($P=1$ to $P=8$) outperforms both linear and fully nonlinear models by reusing structures tasks. For medium data (50 samples per task), more nonlinearity provides additional benefit, while sparse nonlinearity ($P=4$ to $P=16$) performs best. With abundant data (500 samples per task, right), this advantage diminishes as fully nonlinear models can afford task-specific solutions, though sparsely nonlinear architectures remain competitive. (b) Task similarity via Jensen-Shannon divergence between bitcode distributions (mean across 5 independent training runs). Sparse nonlinearity ($P=2,4$) reveals interpretable structure where related tasks share subregions—for instance, pro variants cluster together while anti variants form a separate group. As nonlinearity increases, this structure vanishes as the model has sufficient capacity to learn essentially independent representations for each task, reducing overlap and eliminating the interpretable similarity structure visible at lower $P$.
A natural extension of these results is to examine multi-task learning, where networks must balance task-specific specialization with cross-task generalization. This setting is particularly relevant from a neuroscientific perspective, as biological neural circuits routinely solve multiple related tasks with shared neural resources. We trained AL-RNNs on a suite of 11 cognitive benchmark tasks adapted from Driscoll et al. (2024), presented in randomly interleaved trials (see Appx. A.3 for details). The suite includes pro- and anti-response variants of delayed response, reaction time, and category decision tasks (requiring context-dependent mirror-image mappings), match and non-match variants of delayed match-to-sample, two context-integration tasks that demand selective attention to one of two sensory modalities, and a go/nogo task probing simple threshold detection. These tasks share a common temporal structure and input format, but require distinct computations, making them a natural testbed for understanding how nonlinearity supports flexible reuse of computations across tasks. Fig. 5ashows test accuracy across the 11 tasks as a function of the number of PWL units $P$ and number of task samples used for training. While a significant portion of the multi-task suite requires nonlinearity, with linear models ($P=0$) solving only two of eleven tasks, a single PWL unit ($P=1$) achieves significant performance gains across nearly all tasks simultaneously. The pro/anti task pairs make this particularly transparent. Each variant is individually solvable by a linear system (e.g., mapping a stimulus to the same vs. opposite response), but a single linear network cannot implement both mappings at once, because it would need to "flip" the input-output relationship based on context. Introducing just one nonlinear unit provides this context-dependent switch: it allows the network to route pro and anti computations into different linear regions and thus solve both tasks simultaneously. Crucially, the same single PWL boundary is leveraged to achieve this separation for multiple different tasks simultaneously (see Fig. 24 for bitcodes and Fig. 25 for example latent trajectories). The advantage of sparse nonlinearity depends strongly on the training data. With very limited training data (20 samples per task), sparse nonlinearity ($P=1\text{ to }8$) substantially outperforms both linear and fully nonlinear architectures across most tasks. At intermediate data levels (50 samples per task), sparse models maintain strong performance while fully nonlinear models begin to catch up. With abundant data (500 samples per task), fully nonlinear models achieve the best performance as they can fully converge to task-specific solutions without the constraint of shared subregions. Even in this high-data regime, however, sparse models ($P=2\text{ to }8$) remain competitive. This pattern aligns with the hypothesis that sparse nonlinearity is particularly advantageous in the low-data regime by forcing the network to reuse a small set of nonlinear computations across tasks. To test this "reuse" hypothesis, we examined the network’s distribution of bitcodes on the test set, and quantified similarity via Jensen-Shannon divergence (Fig.  5b, see Fig. 24 for examples of bitcodes). The resulting similarity matrices reveal how tasks share nonlinear resources: at $P=1$, the structure is relatively stark with binary either-or distributions, while $P=4$ and $P=8$ feature more nuanced relationships, where pro variants cluster together, anti variants form a separate group, and complex tasks like match-to-sample develop distinct signatures. This shared structure helps explain the sample efficiency advantage by enforcing cross-task generalization when data is scarce. At full nonlinearity ($P=64$), this structure disappears to a near-diagonal matrix as tasks are spread across independent subregions without recurring shared patterns.